# Building Information Modelling MRR modular reconfigurable robot

### v1

@@ IV-C Task representation
We obtain a formalized task representation by parsing the BIM model and user input as described in Sec. III. The representation of goals, obstacles, tolerances, and robots follows the modeling language introduced by the CoBRA benchmark [45]. A formal definition for our specific use case is included in -B and -C.
@@ IV-D Fitness functions
For motion planning, we use the toolbox for industrial modular robotics (Timor) [44]. We obtain base movement ${\bm{b}}(t)$ and joint angles ${\bm{q}}\left(t\right)$ over time from a black-box path planner
@@
The planner guarantees satisfaction for all constraints ${\mathbb{C}}$ considered in this work, such as collision avoidance. However, inverse kinematics solutions to reach all goals ${\mathbb{G}}$ might not always exist. Further, as we rely on numeric optimization, we might not be able to find them with perfect accuracy. To obtain a scalar expression of the errors ${\mathbf{e}}{}=\left({\bm{t}_{e}},{\bm{R}_{e}}\right)$ (21) given by an imperfect solution, we introduce weighting factors $w_{t},w_{R}$ and error measures $d_{t}:\mathbb{R}^{3}\to\mathbb{R}$ and $d_{R}:SO(3)\to\mathbb{R}$ and define the distance function
@@
While the measures can be chosen arbitrarily, we usually use the Euclidean norm for $d_{t}$ and define $d_{R}$ as the Euler angle of rotation (22). End effector poses can easily be computed using the forward kinematics $FK:\mathbb{R}^{3}\times{\mathbb{Q}}\to SE(3)$. This allows us to obtain the minimum pose error for goal ${\mathbf{g}}_{i}$:
${\mathbf{e}}^{*}_{i}(\bm{\xi}\left(t\right),{\mathbf{p}}_{i})$ $={\mathbf{e}}{({\mathbf{p}}_{i},FK(\bm{\xi}\left(\hat{t}\right)))}\text{, where}$ (7)
$\hat{t}$ $=\operatorname*{arg\,min}_{t}d({\mathbf{e}}({\mathbf{p}}_{i},FK(\bm{\xi}\left(t\right))))\,.$
For better readability, we omit the arguments when they are clear from context and write ${\mathbf{e}}_{i}^{*}$. If the goal error is within the goal tolerance for every goal in a task, we say that the path $\bm{\xi}\left(t\right)$ is feasible.
Let ${\mathbb{O}}_{{\mathbf{m}},i}\left({\bm{q}}\right)$ be the space that is occupied by the $i$th link of robot ${\mathbf{m}}$ with $n$ joints and joint configuration ${\bm{q}}$. The following constraints for joint limits (8), self-collisions (9), and collisions with the environment (10) all evaluate to true or false:
$c_{\text{lim}})$ $\forall t\in[T]:{\bm{q}}\left(t\right)\in{\mathbb{Q}}$ (8)
$c_{\text{sc}})$ $\forall\left(i,t\right)\in\left[n\right]\times[T],\,\forall j<i:$
@@
$c_{\text{col}})$ $\forall\left(i,t\right)\in\left[n\right]\times[T]:{\mathbb{O}}\cap{\mathbb{O}}_{{\mathbf{m}},i}\left(\bm{\xi}\left(t\right)\right)=\emptyset\,.$ (10)
We define the fitness functions for constraint satisfaction of robot ${\mathbf{m}}$ as follows.
@@
${c_{2}}$ $=\min\left(0,\min_{i}\left(\mathrm{count}\left({\mathcal{M}},{\textnormal{m}}_{i}\right)-\mathrm{count}\left({\mathbf{m}},{\textnormal{m}}_{i}\right)\right)\right)$ (12)
@@
${c_{4}}$ =$=\begin{cases}0&\text{, if }\mathrm{feasible}(\bm{\xi}\left({\mathbf{m}},{\mathbb{G}},\{c_{\text{lim}}\},{\mathbb{G}},t\right))\\ d({\mathbf{e}}^{*}_{i})&\text{, otherwise.}\end{cases}$ (14)
${c_{5}}$ =$=\begin{cases}0&\text{, if }\mathrm{feasible}(\bm{\xi}\left({\mathbf{m}},{\mathbb{G}},\{c_{\text{lim}},c_{\text{sc}}\},{\mathbb{G}}\right),t)\\ d({\mathbf{e}}^{*}_{i})&\text{, otherwise.}\end{cases}$ (15)
${c_{6}}$ =$=\begin{cases}0&\text{, if }\mathrm{feasible}(\bm{\xi}\left({\mathbf{m}},{\mathbb{G}},\{c_{\text{lim}},c_{\text{sc}},c_{\text{col}}\},{\mathbb{G}},t\right))\\ d({\mathbf{e}}^{*}_{i})&\text{, otherwise.}\end{cases}$ (16)
The functions in (11) and (12) can be evaluated without obtaining a kinematic or dynamic model of the robot and indicate whether the individual modules of the robot are large enough to reach all goals and if they are available in situ, respectively. The constraint in (13) forces the base not to move after the visual calibration is performed ($t=T_{cal}$, see also Figure 4). The remaining functions successively evaluate whether the joint path is valid regarding the joint limits (14), self-collisions (15), and collisions with the environment (16).
If all constraints are satisfied, we compute the values for our multi-objective fitness criteria. The compactness
@@
indicates the negative value of the total size of all modules combined. High compactness usually allows faster transportation and a smaller lever for the forces induced by drilling.
The robustness fitness function indicates the maximum positioning error our motion planner can account for online. After visual calibration, we obtain a placement error $\Delta_{b}=(\Delta_{x},\Delta_{y},\Delta_{\theta})^{T}$ between the desired and the actual base placement of the robot. The adjustment function $\mathrm{replan}(\bm{\xi}\left(t\right),\Delta_{b})$ indicates whether there is an adjustment $\bm{\varepsilon}$ that, when applied to the joint path ${\bm{q}}\left(t\right)$, results in the end effector of the robot still following the desired motion. Finally, we define the robustness of a robot ${\mathbf{m}}$ with respect to positioning errors of the mobile base as
$f_{r}=$ $\max\{\delta\in[0,1]:\mathrm{replan}(\bm{\xi}\left({\mathbf{m}},{\mathbb{G}},{\mathbb{C}},t\right),\delta\Delta_{max})\}\,.$ (18)
The maximum displacement $\Delta_{max}$ can be chosen based on the available hardware and prior experience. We choose $\begin{bmatrix}20\text{cm}&20\text{cm}&15\deg\end{bmatrix}^{T}$ for our experiments. The above definition does not discriminate between robustness against positioning errors in the $xy$-plane and orientation errors $\Delta_{\theta}$. If necessary, it can easily be extended to multiple robustness scores, one for every possible error. In practice, we experienced that positioning errors were mainly caused by slipping wheels, rendering an equal consideration of the different error modes plausible.
Lastly, the reconfiguration time fitness indicates the time necessary to reconfigure the current module composition ${\mathbf{m}}$ to match a new, desired configuration ${\mathbf{m}}_{d}$. If the current composition is unknown or non-existent, for example, because the modules are currently disassembled, we set ${\mathbf{m}}$ to the empty tuple, so the reconfiguration fitness measures the initial time for setup. Assume that ${\mathbf{m}}$ and ${\mathbf{m}}_{d}$ share $n_{c}$ common modules at the beginning of their kinematic chain. Then, the reconfiguration time fitness is given by
@@
Here, we assume that removing or adding a module takes a time of $t_{m}$, regardless of the module.
@@ V Experiments
We conducted experiments on a drilling use case in the Fraunhofer Italia ARENA at the NOI Techpark in Bolzano, Italy. The experimental setup is shown in Figure 1. It consists of a movable wall with four blocks (brick, sandstone brick, wood, and insulation material) that is modeled in the used BIM file [^4] https://github.com/FraunhoferItalia/rosbim/tree/humble-dev/rosbim_example_models.
@@ V-A Robot modules
We use the robot modules introduced in [46] and shown in Figure 1A, together with an RB-VOGUI mobile base555 https://robotnik.eu/products/mobile-robots/rb-vogui-en/ with a custom cabinet top setup and two lidars (SICK expert S300). In total, we use one mounting base, three elbow- and straight joint modules each, three passive links666 https://github.com/FraunhoferItalia/fhi_resources with a length of $100$, $200$, and $500$ millimeters, respectively, as well as a custom-made end effector, composed of a drill tool (Einhell Herocco 18/20 with custom 3D printed support) and an RGB-D camera (Intel Realsense D435i).
V-B Validation in simulation
(a) Normalized compactness and robustness of Pareto-optimal results within 100 individual runs. The heights of the boxes indicate the relative frequency of a particular fitness score observed. The blue line indicates the Pareto optimal results concerning all trials.
@@ V-B Validation in simulation
Figure 3a shows the Pareto-optimal solutions of each trial regarding compactness and robustness. The compactness is normalized for better readability, i.e., 00 indicates the largest and $1$ the most compact robot present in the final solution set, respectively. As our optimization is not complete, the theoretic Pareto front remains unknown. Instead, we report the evidence-based Pareto front by aggregating over all trials in blue. Further, Figure 3b shows the tradeoff between robustness and reconfiguration time, starting from an exemplary initial configuration.
@@ V-C Real-world experiments
Figure 4 shows a simplified version of the behavior tree that consists of one sequence of robot actions used in the experiments. The full tree additionally includes fallback nodes for error handling. Pictures of the drilled holes can be seen in the most-right snapshot in Figure 4 and in Figure 6. A full video of the mission execution is available on the project website modular-construction-robot.cps.cit.tum.de.
Figure 5: Among the conducted experiments, we measure absolute errors between $6$ cm and $16$ cm regarding the base position and between $0.7^{\circ}$ and $12.6^{\circ}$ regarding its orientation. The figure shows the positioning errors of the base with respect to the ArUco markers and the maximum joint angle adjustment during execution of the drilling task.
Based on prior experience with the mobile base, we chose a minimum robustness score of $\delta=0.8$. Consecutively, we selected two solutions from our optimization; one with a six-degrees-of-freedom (DoF) robot arm and one with a five-DoF robot arm. For every configuration, we performed three drilling missions on the sandstone brick.
@@
where ${\bm{q}}$ is the initially planned joint configuration, and $\hat{{\bm{q}}}$ is the one observed during execution. We chose this metric as it can easily be compared between robots with varying DoF; however, the overall trend shown in Figure 5 persists over different choices of norms.
The first experiment was conducted with the six-DoF manipulator. Figure 6 shows the sandstone brick after drilling three holes in a horizontal line. The relative vertical precision was within one centimeter.
@@
Initially, the modules were disassembled, as depicted in Figure 1A. Upon receiving the configuration for the robot (a six-DoF setup consisting of the following module sequence: straight, elbow, passive 100mm, straight, elbow, straight, elbow, end effector), the operator assembled the robot in ten minutes. For the second experiment, the robot consisted of the following module sequence: straight, elbow, passive 100mm, straight, elbow, elbow, end effector. The time required to reconfigure the robot and launch the mission was reduced to about five minutes, as the first five modules were shared with the previous configuration.
V-D Discussion

### v2

@@ IV-C Task representation
We obtain a formalized task representation by parsing the BIM model and user input as described in Sec. III. The representation of goals, obstacles, tolerances, and robots follows the modeling language introduced by the CoBRA benchmark [45]. A formal definition for our specific use case is included in Appendices -B and -C.
@@ IV-D Fitness functions
We obtain base movement ${\bm{b}}(t)$ and joint angles ${\bm{q}}\left(t\right)$ over time from a black-box path planner
@@
The planner guarantees satisfaction for all constraints ${\mathbb{C}}$ considered in this work, such as collision avoidance. However, inverse kinematics solutions to reach all goals ${\mathbb{G}}$ may not always exist. Furthermore, since we rely on numeric optimization, we might not be able to find them with perfect accuracy. To obtain a scalar expression of the errors ${\mathbf{e}}{}=\left({\bm{t}_{e}},{\bm{R}_{e}}\right)$ defined in (24), we introduce weighting factors $w_{t},w_{R}$ and error measures $d_{t}:\mathbb{R}^{3}\to\mathbb{R}$ and $d_{R}:SO(3)\to\mathbb{R}$ and define the distance function
@@
While the measures can be chosen arbitrarily, we usually use the Euclidean norm for $d_{t}$ and define $d_{R}$ as the Euler angle of rotation, as detailed in (25). The end-effector poses can be easily computed using forward kinematics $FK:\mathbb{R}^{3}\times{\mathbb{Q}}\to SE(3)$. This allows us to obtain the minimum pose error for goal ${\mathbf{g}}_{i}$:
${\mathbf{e}}^{*}_{i}$ $={\mathbf{e}}{\Bigl{(}{\mathbf{p}}_{i},FK\bigl{(}\bm{\xi}\left(\hat{t}\right)\bigr{)}\Bigr{)}}\text{, where}$ (8)
$\hat{t}$ $=\operatorname*{arg\,min}_{t}d\Bigl{(}{\mathbf{e}}\bigl{(}{\mathbf{p}}_{i},FK(\bm{\xi}\left(t\right))\bigr{)}\Bigr{)}\,.$
If the goal error is within the goal tolerance for every goal in a task, we say that the path $\bm{\xi}\left(t\right)$ is feasible.
Let ${\mathbb{O}}_{{\mathbf{m}},i}\left({\bm{q}}\right)$ be the space occupied by the $i$th link of robot ${\mathbf{m}}$ with joint configuration ${\bm{q}}$. The following constraints for joint limits (9), self-collisions (10), collisions with the environment (11), and torque limits (12), are later used for the constraint functions $c_{j}$ and all evaluate to true or false:
$c_{\text{lim}}$ $\iff\forall t\in[T]:{\bm{q}}\left(t\right)\in{\mathbb{Q}}$ (9)
$c_{\text{sc}}$ $\iff\forall\left(i,t\right)\in\left[n\right]\times[T],\,\forall j<i:$
@@
$c_{\text{col}}$ $\iff\forall\left(i,t\right)\in\left[n\right]\times[T]:{\mathbb{O}}\cap{\mathbb{O}}_{{\mathbf{m}},i}\left(\bm{\xi}\left(t\right)\right)=\emptyset$ (11)
$c_{\tau}$ $\iff\forall t\in[T]:\bm{\tau}(t,{\bm{f}}_{ext}(t))\in[\underline{\bm{\tau}},\overline{\bm{\tau}}]\,.$ (12)
Here, $\bm{\tau}(t,\bm{f}_{ext}(t))$ denotes robot joint torques under consideration of an external payload ${\bm{f}}_{ext}(t)$. We define the fitness functions for the constraint satisfaction of the robot ${\mathbf{m}}$ as follows.
@@
${c_{2}}$ $=\min\left(0,\min_{i}\left(\mathrm{count}\left({\textnormal{m}}_{i},{\mathcal{M}}\right)-\mathrm{count}\left({\textnormal{m}}_{i},{\mathbf{m}}\right)\right)\right)$ (14)
@@
${c_{4}}$ =$=\begin{cases}0&\text{, if }\mathrm{feasible}(\bm{\xi}\left({\mathbf{m}},{\mathbb{G}},\{c_{\text{lim}}\},{\mathbb{G}},t\right))\\ -d({\mathbf{e}}^{*}_{i})&\text{, otherwise.}\end{cases}$ (16)
${c_{5}}$ =$=\begin{cases}0&\text{, if }\mathrm{feasible}(\bm{\xi}\left({\mathbf{m}},{\mathbb{G}},\{c_{\text{lim}},c_{\text{sc}}\},{\mathbb{G}}\right),t)\\ -d({\mathbf{e}}^{*}_{i})&\text{, otherwise.}\end{cases}$ (17)
${c_{6}}$ =$=\begin{cases}0&\text{, if }\mathrm{feasible}(\bm{\xi}\left({\mathbf{m}},{\mathbb{G}},\{c_{\text{lim}},c_{\text{sc}},c_{\text{col}}\},{\mathbb{G}},t\right))\\ -d({\mathbf{e}}^{*}_{i})&\text{, otherwise.}\end{cases}$ (18)
$c_{7}$ $=\sum_{t=1}^{T}\sum_{i=1}^{n}\min\left(0,\overline{\tau}_{i}-\tau_{i}(t,{\bm{f}}_{ext}),\tau_{i}(t,{\bm{f}}_{ext})-\underline{\tau}_{i}\right)$ (19)
The functions in (13) and (14) can be evaluated without obtaining a kinematic or dynamic model of the robot and indicate whether the individual modules of the robot are large enough to reach all goals and if they are available on-site, respectively. The fitness function in (15) forces the base not to move after visual calibration is performed ($t=T_{cal}$, see also Figure 4). The next three functions successively assess whether the joint path is valid with respect to joint limits (16), self-collisions (17), and collisions with the environment (18). Finally, (19) is zero if and only if all torque limits are kept.
If all constraints are satisfied, we compute the values for our multiobjective fitness criteria. The compactness
@@
indicates the negative value of the total size of all modules combined. High compactness usually allows for faster transport and a shorter lever for the forces induced by drilling.
The robustness fitness function indicates the maximum placement error $\Delta_{b}=(\Delta_{x},\Delta_{y},\Delta_{\theta})^{T}$ our motion planner can account for online. The adjustment function $\mathrm{replan}(\bm{\xi}\left(t\right),\Delta_{b})$ indicates whether there is an adjustment that when applied to the joint path ${\bm{q}}\left(t\right)$, results in the end effector of the robot still following the desired motion. Finally, we define the robustness of a robot ${\mathbf{m}}$ with respect to positioning errors of the mobile base as
[…]
$\quad\mathrm{replan}(\bm{\xi}\left({\mathbf{m}},{\mathbb{G}},{\mathbb{C}},t\right),\delta\Delta_{max})=\mathtt{true}\}\,.$
The maximum displacement $\Delta_{max}$ can be chosen based on the available hardware and prior experience. We chose $\begin{bmatrix}20\text{cm}&20\text{cm}&15\deg\end{bmatrix}^{T}$ for our experiments. The above definition does not discriminate between robustness against positioning errors in the $xy$-plane and orientation errors $\Delta_{\theta}$. If necessary, the robustness score can be easily extended to multiple scores, one for every possible error.
Lastly, the reconfiguration time fitness indicates the time necessary to reconfigure the current module composition ${\mathbf{m}}$ to match a new desired configuration ${\mathbf{m}}_{d}$. If the current composition is unknown or nonexistent, for example, because the modules are currently disassembled, we set ${\mathbf{m}}$ to the empty tuple, so the reconfiguration fitness measures the initial time for setup. Assume that ${\mathbf{m}}$ and ${\mathbf{m}}_{d}$ share $n_{c}$ common modules at the beginning of their kinematic chain. Then, the reconfiguration time fitness is chosen as
@@
Here, we assume that removing or adding a module takes time $t_{m}$, regardless of the module.
@@ V Experiments
We conducted experiments on a drilling use case in the Fraunhofer Italia ARENA at the NOI Techpark in Bolzano, Italy. The experimental setup is shown in Figure 1. It consists of a movable wall with four blocks (brick, sandstone brick, wood, and insulation material) that is modeled in the used BIM file333 https://github.com/FraunhoferItalia/rosbim/tree/humble-dev/rosbim_example_models .
@@ V-A Robot modules
We use the robot modules introduced in [46] and shown in Figure 1A, together with an RB-VOGUI mobile base444 https://robotnik.eu/products/mobile-robots/rb-vogui-en/ with a custom cabinet top setup and two lidars (SICK expert S300). In total, we use one mounting base, three elbow- and straight joint modules each, three passive links555 https://github.com/FraunhoferItalia/fhi_resources with a length of $100$, $200$, and $500$ millimeters, respectively, as well as a custom-made end effector composed of a drill tool (Einhell Herocco 18/20 with custom 3D printed support) and an RGB-D camera (Intel Realsense D435i).
V-B Drilling in simulation
(a) Normalized compactness and robustness of Pareto-optimal results within 100 individual runs. The heights of the boxes indicate the relative frequency of a particular fitness score observed. The green line indicates the Pareto optimal results for all trials.
@@ V-B Drilling in simulation
Each of the 100 runs resulted in at least eleven valid solutions, all of which had five or six joints. On average, the Pareto front consisted of $3.7$ distinct individuals. For $99\%$ of the Pareto-optimal solutions, the first two joints were a straight, then an elbow joint. Performance differences resulted from the remaining modules, which were distributed in various ways. Interestingly, we observed a strong positive correlation between solution robustness and using the longest passive link (Pearson correlation coefficient $\rho=0.54,p<0.01$). The robustness of the calibration was also correlated with the use of the third straight joint ($\rho=0.32,p<0.01$). In contrast, using the third elbow joint was positively correlated with the compactness of Pareto-optimal solutions ($\rho=0.52,p<0.01$). Figure 3a shows the Pareto-optimal solutions of each trial regarding compactness and robustness. The compactness is normalized for better readability; that is, 00 indicates the largest robot and $1$ the most compact robot present in the final solution set, respectively. As our optimization is not complete, the theoretic Pareto front remains unknown. Instead, we report the evidence-based Pareto front by aggregating all trials. Further, Figure 3b shows the trade-off between robustness and reconfiguration time, starting from an exemplary initial configuration.
In addition to our algorithm, we performed ten optimizations using the lexicographic genetic approach introduced in [36]. As this algorithm does not support multiple objectives, we optimized for compactness only. Other fitness functions, constraints, and hyperparameters remained unchanged. Normalized by the results of our algorithm, we observed an average compactness score of $0.85$ (maximum: $1.0$) and an average robustness score of $0.04$ (maximum: $0.1$).
@@ V-C Real-world experiments
To evaluate the real-world performance of our system, we conducted a drilling experiment covering the entire workflow from task specification to execution. Figure 4 shows a simplified version of the behavior tree that consists of one sequence of robot actions used in the experiments. The full tree also includes fallback nodes for error handling. The drilling task includes multiple stages. First, the mobile robot base navigates towards the experimental setup, using waypoints identified from the BIM model. The robot arm then unfolds to a calibration pose, and the base position is refined based on the ArUco markers, scanned by a camera mounted on the end effector. After calibration, the initial planned robot trajectory is adapted online to counteract the final positioning error of the mobile base. The robustness of the robot plays a crucial role in this step. If the initial trajectory is close to the joint limits of the robot, a large positioning error might lead to failure at this stage, requiring recalibration. Finally, the robot performs the approach and drill motions, where the trajectory is executed using the impedance controller of the robot arm. Pictures of the drilled holes can be seen in the most-right snapshot in Figure 4 and in Figure 6. A complete video of the mission execution is available on the project website modular-construction-robot.cps.cit.tum.de.
Figure 5: Among the conducted experiments, we measure absolute errors between $6$ cm and $16$ cm regarding the base position and between $0.7^{\circ}$ and $12.6^{\circ}$ regarding its orientation. The figure shows the positioning errors of the base with respect to the ArUco markers and the maximum joint angle adjustment during the execution of the drilling task.
Based on previous experience with the mobile base, we chose a minimum robustness score of $\delta=0.8$. Consecutively, we selected two solutions from our optimization: one with a six-degree-of-freedom (DoF) robot arm and one with a five-DoF robot arm. For each configuration, we performed three drilling missions on the sandstone brick.
@@
where ${\bm{q}}$ is the initially planned joint configuration, and $\hat{{\bm{q}}}$ is the one observed during execution. We chose this metric as it can easily be compared between robots with varying DoF; however, the general trend shown in Figure 5 persists over different choices of norms.
@@
Figure 6 shows the sandstone brick after drilling three holes in a horizontal line. The relative vertical precision was within one centimeter. Initially, the modules were disassembled, as depicted in Figure 1A. Upon receiving the configuration for the first robot (a six-DoF setup consisting of the following module sequence: straight, elbow, passive 100mm, straight, elbow, straight, elbow, end effector), the operator assembled the robot in ten minutes. The second robot consisted of the following module sequence: straight, elbow, passive 100mm, straight, elbow, elbow, end effector. The time required to reconfigure the robot and launch the mission was reduced to about five minutes, since the first five modules were shared with the previous configuration.
V-D Spray painting in simulation
Figure 7: Simulation of spray painting: The thin red line shows the target trajectory.
Figure 8: Normalized compactness and robustness of Pareto-optimal results for spray painting within 100 individual runs. The heights of the boxes indicate the relative frequency of a particular fitness score observed. The green line indicates the Pareto optimal results for all trials.
Our framework can be adapted to various construction tasks, as shown by the following example for spray painting. We adapt the BIM model of the drilling task by including an additional wall. Additionally, we replaced the end-effector module with a custom end effector for spray painting, first introduced in [15]. In a slightly modified user interface, users can select a rectangular area to be painted instead of drilling positions, for which a trajectory covering this area is automatically computed. We also adjust the constraints by removing the drilling payload and allowing the robot base to move during task execution to cover larger areas. Similarly to drilling, task performance is independent of the end-effector orientation around its local z-axis, so the goal tolerances need not be adjusted. Other than in the drilling task, the main challenge in spray painting is not the accurate positioning of the end effector. Instead, reducing the number of recalibrations triggered by base movement is important. Consequently, we use the negative number of required recalibrations as the robustness score for this task.
We performed 100 optimization trials, 99 of which resulted in at least one valid robot morphology. On average, the Pareto front consisted of 3.1 distinct individuals. In contrast to the drilling task, 18% of the robots had a passive link as their first module, as shown in Figure 7. The robots for spray painting consisted of an average of $8.2$ modules, compared to an average of $6.5$ for drilling. Figure 8 shows the normalized scores for Pareto-optimal solutions for spray painting. Similarly to the drilling task, we observed a trade-off between the compactness and robustness of the solutions. Again, this was also reflected in a positive correlation between the use of the longest passive link and the robustness $(\rho=0.41,p<0.01)$. All valid solutions used the three elbow joints available.
V-E Discussion
The results show the trade-offs between compactness, robustness, and configuration time. As shown by the results for our baseline algorithm, single-objective optimization approaches tend to provide a single solution on the Pareto front only. Theoretically, the full Pareto front can still be recovered using single-objective optimization. However, this would require carefully chosen weighting factors and scale exponentially with the number of objectives to be considered. Our results show that explicit knowledge about the trade-off between competing objectives is crucial for practical applications. The Pareto front shown in Figure 3a indicates that a good robustness score of $0.4$ can already be achieved with a robot almost as compact as the most compact candidate. This relation does not hold the other way around: The most robust robots are significantly less compact than those with robustness scores of $0.9$ or less.